# Narratives of Collective Action in YouTube’s Discourse on Veganism

### v1

@@ Results
m: 58
n: 76
v: 131
m: 61
n: 75
@@
Narratives - communal-oriented
Narratives - agency-oriented
@@
Code snippets provided by Google Developers for the use of YouTube API are licensed under the Apache 2.0 License. Python packages used to process YouTube data, namely Youtube Transcript, langdetect and Whisper are licensed under an MIT license, while Llama-2-70B-chat model employed in the mapping validation is an open-source LLM released under a commercial use license [^1] https://ai.meta.com/llama/license/. Mformer models used in the extraction of moral scores are shared on HuggingFace and licensed under an MIT License, while S-BERT model for the extraction of sentence embeddings is licensed under the Apache License 2.0. The collective action dictionary by (Smith, McGarty, and Thomas 2018) is publicly available.

### v2

@@ Acknowledgments
m: 56
n: 147
v: 125
m: 49
n: 88
Figure 7 outlines the distribution of videos by mapped narrative type.
Figure 7: Number of videos by narrative type, according to the reference moral vegan ideology framework.
@@
Narratives - communal-oriented Narratives - agency-oriented
@@
Code snippets provided by Google Developers for the use of YouTube API are licensed under the Apache 2.0 License. Python packages used to process YouTube data, namely Youtube Transcript, langdetect and Whisper are licensed under an MIT license, while Llama-2-70B-chat model employed in the mapping validation is an open-source LLM released under a commercial use license [^2] https://ai.meta.com/llama/license/. Mformer models used in the extraction of moral scores are shared on HuggingFace and licensed under an MIT License, while S-BERT model for the extraction of sentence embeddings is licensed under the Apache License 2.0. The collective action dictionary by Smith, McGarty, and Thomas (2018) is publicly available.